# Quantum Oppenheimer-Snyder models in loop quantum cosmology with Lorentz term

### v1

@@ IV THERMODYNAMIC PROPERTIES
In the preceding sections, we derived the metric for a spacetime without a cosmological constant and analyzed its properties. However, our Universe is undergoing an accelerated expanding phase. This leads us to investigate the effect of the cosmological constant on spacetime, as it introduces important modifications to the structure and dynamics of spacetime. Moreover, by introducing a negative cosmological constant in black hole physics will bring us richer physics such as Hawking-Page phase transition in Schwarzschild-AdS case. Hence, the inclusion of a cosmological constant in a black hole solution have significance both in the practical and the theoretical sense. In order to investigated the thermodynamic properties of the quantum black holes in the extended formalism, one usually need to extend the effective metric to include a negative cosmological constant [58, 59]. By substituting the previous Friedmann equation Eq. (2) with the modified Friedmann equation [60] that incorporates the cosmological constant, and following the same derivation process as in the preceding sections, we can derive the metric that includes the cosmological constant, which is
@@
$F(r)=1-\frac{r^{2}}{\gamma^{2}\Delta}f(\rho)\Big{(}1-f(\rho)\Big{)}(1-\frac{\rho}{\rho_{c}}),\qquad f(\rho)=\frac{1\pm\sqrt{1-\frac{\rho}{\rho_{c}}}}{2(1+\gamma^{2})},\qquad\rho=\frac{M}{\frac{4}{3}\pi r^{3}}+\frac{\Lambda}{8\pi G}.$
As discussed earlier, there exist two distinct branches of solutions. The $f(\rho)_{-}$ branch corresponds to a geometry that asymptotically approaches a flat spacetime, while the $f(\rho)_{+}$ branch asymptotically approaches a de Sitter spacetime with an effective cosmological constant [21, 61]. Since the bounce is expected to occur in both time directions, two scenarios are possible: a transition from the $f(\rho)_{-}$ branch to the $f(\rho)_{+}$ branch, or vice versa. However, the transition from the $f(\rho)_{+}$ branch to the $f(\rho)_{-}$ branch appears to be more physically plausible. The de Sitter–like asymptotic features of the $f(\rho)_{+}$ branch motivates further investigation into its thermodynamic properties. Therefore, we investigate the thermodynamic properties of the $f(\rho)_{+}$ branch in detail.
@@ IV.1 Anti-de Sitter spacetime
Due to the rich and intricate thermodynamic properties exhibited by AdS spacetimes, this subsection will primarily focus on the $f(\rho)_{-}$ branch has a negative cosmological constant. In particular, we will discuss the thermodynamic properties in this context, including the Hawking temperature, heat capacity, and entropy, which are central to understanding the behavior of black holes and phase transitions in AdS backgrounds.
@@
where $\mathcal{K}$ is the surface gravity of the black hole, $r_{h}$ is the event horizon radius, which is the largest root of the Eq. (23) with the $f(\rho)_{-}$ branch. In Fig. 15, we compare the Hawking temperature for two types of black holes: the LQG black hole and the classical counterpart. As shown in the graph, the temperature continuously decreases as r increases for the classical scenario, and the temperature diverges, tending towards infinity. In contrast, the LQG black hole exhibits a different behavior. Initially, the temperature increases as r grows, reaching a maximum value. Afterward, the temperature decreases, eventually approaching the same behavior as the classical black hole at large r. In some modified gravity theories, the Hawking temperature exhibits a similar behavior [58, 62].
@@
From the Eq.(23) with the $f(\rho)_{-}$ branch, we can use the event horizon $r_{h}$ to express the black hole’s mass. In Fig.15, we plot the curve of the heat capacity for $r_{h}$ and compare it with the classical scenario. As shown in the graph, we can find that the LQG black hole compared with the classical scenario has an extra phase transition. Similar phenomena can be observed in some modified gravity [58].
@@
In Fig. 16, we display the behavior of the ratio $\frac{S}{r_{h}}$ as a function of the horizon radius $r_{h}$ and compare the LQG black hole with the classical scenario. As shown in the graph, we can see that the ratio $\frac{S}{r_{h}}$ for the LQG black hole exhibits a non-monotonic behavior: it initially increases with $r_{h}$, then decreases, and eventually approaches the classical result at large $r_{h}$.
In general, the entropy of a classical black hole satisfy the area law, $S=\frac{A}{4}$, where $A$ is the area of the event horizon. However, in cases involving quantum corrections and higher-dimensional gravity, the area law no longer holds. Compared to the general case, an additional logarithmic correction term appears [67, 68, 58]. In our work, the entropy also contains a logarithmic correction term, which is:
@@
Figure 16: The ratio $\frac{S}{r_{h}}$ of the LQG and Schwarzschild black holes as a function of even horizon $r_{h}$. Parameters used: $\Lambda=-0.001$.
@@
The epoch of $f(\rho)_{+}$ branch does not have a cosmological constant, but its quantum geometric effects are equivalent to a positive effective cosmological constant of Planckian order [21]. The effective cosmological constant is $\Lambda=\frac{3}{(1+\gamma^{2})^{2}\Delta}$, resulting in an de-Sitter epoch. To investigate how quantum geometric effects influence black hole thermodynamics, we compare the black hole solution in the $f(\rho)_{+}$ branch with the dS Schwaschild black hole, which has the same effective cosmological constant $\Lambda$. For the $f(\rho)_{+}$ branch of the Eq. (23), we define the outer horizon as the event horizon $r_{h}$.
In Fig. 18, we investigate the Hawking temperature how to effect of the event horizon. In the dS Schwaschild black hole, the Hawking temperature increases linearly with $r_{h}$. In contrast, for the LQG black hole, the Hawking temperature first decreases with increasing $r_{h}$, and then increases linearly. Moreover, the temperature increases at different rates in the two cases, indicating that although the temperature profiles have a similar overall shape, their thermodynamic behaviors may not be identical. It should be noted that the Hawking temperature is usually positive, therefore, the temperature of the classical black hole starts from zero.
We also investigate the relationship of the heat capture with the even horizon $r_{h}$ on the Fig. 18. We can see that, compared to the Schwaschild-dS black hole, the LQG black hole has an additional phase transition, which reflects the influence of quantum effects on its thermodynamic properties. At large horizon radius $r_{h}$, the heat capacity of both the Schwaschild-dS black hole and the LQG black hole becomes negative, indicating that both of them are thermodynamically unstable.
Figure 17: Hawking temperature of the LQG and the classical Schwarzschild-dS black holes as a function of event horizon $r_{h}$.
Figure 18: Heat capacity of the LQG and the classical Schwarzschild-dS black holes as a function of event horizon $r_{h}$.
@@ V CONCLUSIONS AND DISCUSSIONS
This study systematically explores the qOS framework within Loop Quantum Cosmology (LQC) augmented by a Lorentz term, establishing its efficacy in characterizing quantum-gravitational dynamics. The model exhibits an asymmetric quantum bounce bifurcated into two phases: the $f(\rho)_{+}$ branch governing de Sitter epoch, and the $f(\rho)_{-}$ branch describing our classical post-bounce universe. Our investigation concentrates on the phenomenologically relevant $f(\rho)_{-}$ sector.
We systematically investigate how angular momentum $l$ and the mass $M$ influence on effective potential, QNMs, and the late-time tail under scalar perturbations. We find that when angular momentum $l$ increases, the peak of the effective potential increases, and the quasinormal frequencies become longer-lived and dominate the time-domain signal. When the mass of black hole $M$ increases, the peak of the effective potential decreases. The late-time tails of black holes with different masses are similar and follow an inverse power-law behavior. To gain a better understanding of the differences in the structure of spacetime for the Schwarzschild black hole, the black hole in the qOS model, and the black hole in the qOS model with Lorentz term, we compared their effective potential and found a small deformation near the even horizon. Furthermore, we compared their time-domain profiles. The profile of the black hole in the qOS model closely matches that of the Schwarzschild black hole, while the black hole in the qOS model with Lorentz term exhibits a slower exponential decay rate. All three configurations display universal inverse power-law tails behavior across all mass configurations.
@@
Note that our universe is undergoing accelerated expansion, we consider the $f(\rho)_{-}$ branch epoch having the cosmological constant. At the same time, the AdS black hole has rich thermodynamic properties, so we mainly investigate the $f(\rho)_{-}$ branch that has a negative cosmological constant. For small $r_{h}$, unlike the classical case where the Hawking temperature decreases monotonically from infinity, the temperature in the LQG black hole initially increases from zero, reaches a maximum, and then decreases. As $r_{h}$ becomes large, the behavior of the LQG black hole becomes similar to the classical scenario. For the heat capacity, we find that the LQG black hole exhibits an additional phase transition at a small $r_{h}$ compared to the classical scenario. When $r_{h}$ becomes large, the behavior of the LQG black hole coincides with that of the classical case. We also investigate the entropy of the LQG black holes. Compared to the classical case, the entropy-to-radius ratio $\frac{S}{r_{h}}$ of the LQG black hole initially increases with $r_{h}$, reaches a maximum, and then decreases, eventually converging to the classical case at large $r_{h}$. This indicates that besides the classical area law $S=\frac{A}{4}$ term, the LQG black hole also has an additional logarithmic correction term due to the quantum gravity effect.
Finally, since the $f(\rho)_{+}$ branch asymptotically approaches a de Sitter spacetime, we also investigate how quantum effects influence on it’s thermodynamic properties. We find that for the LQG black hole, the Hawking temperature first decreases with increasing horizon radius $r_{h}$, and then increases linearly, while the Hawking temperature of Schwaschild-dS black hole monotonically increases with $r_{h}$. However, the rate of increase differs between the two cases. The heat capacity of the LQG black hole exhibits an additional phase transition at small $r_{h}$ compared to the Schwaschild-dS black hole. And at large horizon radius $r_{h}$, the heat capacity of the LQG black hole becomes negative, indicating that both black holes are thermodynamically unstable.

### v2

@@ IV THERMODYNAMIC PROPERTIES
In the preceding sections, we derived the metric corresponding to the post-bounce branch of the spacetime without a cosmological constant for both the cosmological and collapse scenarios, and analyzed its properties. As discussed earlier, the metric of the spacetime on the post-bounce branch differs between the cosmological and collapse cases. In the cosmological case, the vacuum exterior metric corresponding to the post-bounce branch is denoted by $f(\rho)_{-}$, whose geometry asymptotically approaches flat spacetime. In contrast, for the collapse case, the vacuum exterior metric corresponding to the post-bounce branch is given by $f(\rho)_{+}$, which asymptotically approaches a de Sitter spacetime with an effective cosmological constant [23, 29].
However, our Universe is currently undergoing an accelerated expansion, the simplest possible explanation of which is the presence of a positive cosmological constant ($\Lambda>0$). This motivates us to investigate the effect of the cosmological constant on the spacetime, as it introduces important modifications to its structure and dynamics. Moreover, the success of the AdS/CFT correspondence has drawn considerable attention to the case with a negative cosmological constant ($\Lambda<0$). Introducing a negative cosmological constant in black hole physics also leads to richer phenomena. Hence, the inclusion of a cosmological constant in the solution have significance both in the practical and the theoretical sense. To investigate the thermodynamic properties of the vacuum exterior metrics for the $f(\rho)_{-}$ and $f(\rho)_{+}$ branches, it is necessary to extend the metrics to include the cosmological constant term [66, 67]. By substituting the previous Friedmann equation Eq. (2) with the modified Friedmann equation [68] that incorporates the cosmological constant, and following the same derivation process as in the preceding sections, we can derive the metric that includes the cosmological constant, which is
@@
$F(r)=1-\frac{r^{2}}{\gamma^{2}\Delta}f(\rho)\Big(1-f(\rho)\Big)(1-\frac{\rho}{\rho_{c}}),\qquad f(\rho)_{\pm}=\frac{1\pm\sqrt{1-\frac{\rho}{\rho_{c}}}}{2(1+\gamma^{2})},\qquad\rho=\frac{M}{\frac{4}{3}\pi r^{3}}+\frac{\Lambda}{8\pi G}.$
To capture the richer thermodynamic features of the black hole, we incorporate a negative cosmological constant into the vacuum exterior metric of the $f(\rho)_{-}$ branch. As discussed earlier, the metric of the $f(\rho)_{+}$ branch does not contain a cosmological constant, but its quantum geometric effects are equivalent to a positive effective cosmological constant, and resulting in an de-Sitter epoch. To examine how the quantum geometric effects influence the thermodynamic properties, we compare the thermodynamic behavior of this spacetime with that of the classical case. Furthermore, as we show below, incorporating a negative cosmological constant into the metric of this case does not alter our main conclusions.
@@ IV.1 Anti-de Sitter spacetime
Due to the rich and intricate thermodynamic properties exhibited by AdS spacetimes, this subsection will primarily focus on the metric of the $f(\rho)_{-}$ branch with a negative cosmological constant. In particular, we will discuss the thermodynamic properties in this context, including the Hawking temperature, heat capacity, and entropy, which are central to understanding the behavior of black holes and phase transitions in AdS backgrounds.
@@
where $\mathcal{K}$ is the surface gravity of the black hole, $r_{h}$ is the event horizon radius, which is the largest root of the Eq. (35) with the $f(\rho)_{-}$ branch. In Fig. 15, we compare the Hawking temperature for two types of black holes: the LQG-inspired black hole and the classical counterpart. As shown in the graph, the temperature monotonically decreases as $r_{h}$ increases in the classical scenario. When $r_{h}$ approaches zero, the temperature diverges, tending towards infinity. In contrast, the LQG-inspired black holes exhibits a different behavior. The temperature starts at the extremal case, where the $f(\rho)_{-}$ branch of the metric in Eq. (35) has a single horizon. Initially, the temperature increases as $r_{h}$ grows, reaching a maximum value. Afterward, the temperature decreases, eventually approaching the same behavior as the classical black hole at large $r_{h}$. In some modified gravity theories, the Hawking temperature exhibits a similar behavior [66, 69].
@@
From the Eq.(35) with the $f(\rho)_{-}$ branch, we can use the event horizon $r_{h}$ to express the black hole’s mass. In Fig.15, we plot the curve of the heat capacity for $r_{h}$ and compare it with the classical scenario. As shown in the graph, we can find that the LQG-inspired black hole compared with the classical scenario has an extra phase transition. Similar phenomena can be observed in some modified gravity [66].
@@
In Fig. 16, we display the behavior of the ratio $\frac{S}{A}$ as a function of the horizon radius $r_{h}$ and compare the LQG-inspired black hole with the classical scenario, where $A$ is the area of the event horizon. As shown in the graph, we can see that the ratio $\frac{S}{A}$ for the LQG-inspired black hole exhibits a non-monotonic behavior: it initially increases with $r_{h}$, then decreases, and eventually approaches the classical result at large $r_{h}$.
In general, the entropy of a classical black hole satisfy the area law, $S=\frac{A}{4}$. However, in cases involving quantum corrections and higher-dimensional gravity, the area law no longer holds. Compared to the general case, an additional logarithmic correction term appears [74, 75, 66]. In our work, the entropy also contains a logarithmic correction term, which is:
@@
Figure 16: The ratio $\frac{S}{A}$ of the LQG and Schwarzschild black holes as a function of even horizon $r_{h}$. Parameters used: $\Lambda=-0.001$.
@@
The metric of the $f(\rho)_{+}$ branch does not contain a cosmological constant, but its quantum geometric effects are equivalent to a positive effective cosmological constant of Planckian order [23]. The effective cosmological constant is $\Lambda_{eff}=\frac{3}{(1+\gamma^{2})^{2}\Delta}$, resulting in an de-Sitter epoch.
To investigate how quantum geometric effects influence its thermodynamics, we compare this solution with the de Sitter Schwarzschild black hole, which has the same effective cosmological constant $\Lambda_{eff}$. For the $f(\rho)_{+}$ branch of the Eq. (35), we define the cosmological horizon as $r_{c}$.
In Fig. 18, we investigate the Hawking temperature how to effect of the cosmological horizon. In the de Sitter Schwarzschild black hole, the Hawking temperature increases linearly with $r_{c}$. In contrast, for the LQG-inspired black hole, the Hawking temperature first decreases with increasing $r_{c}$, and then increases linearly. Moreover, the temperature increases at different rates in the two cases, indicating that although the temperature profiles have a similar overall shape, their thermodynamic behaviors may not be identical. It should be noted that the Hawking temperature is usually positive, therefore, the temperature of the classical black hole starts from zero.
We also investigate the relationship of the heat capture with the cosmological horizon $r_{c}$ on the Fig. 18. We can see that, compared to the de Sitter Schwarzschild black hole, the LQG solution has an additional phase transition, which reflects the influence of quantum effects on its thermodynamic properties. At large horizon radius $r_{c}$, the heat capacity of both the de Sitter Schwarzschild black hole and the LQG solution becomes negative, indicating that both of them are thermodynamically unstable.
Figure 17: Hawking temperature of the LQG and the classical Schwarzschild-dS black holes as a function of cosmological horizon $r_{c}$.
Figure 18: Heat capacity of the LQG and the classical Schwarzschild-dS black holes as a function of cosmological horizon $r_{c}$.
We can also incorporate a negative cosmological constant into this metric, while ensuring that its absolute value does not exceed the effective cosmological constant $\Lambda_{eff}$. We find that the conclusion remains unchanged.
@@ V CONCLUSIONS AND DISCUSSIONS
This study systematically explores the qOS framework within Loop Quantum Cosmology (LQC) augmented by a Lorentz term, establishing its efficacy in characterizing quantum-gravitational dynamics. The model exhibits an asymmetric quantum bounce bifurcated into two phases: the $f(\rho)_{+}$ branch governing de Sitter epoch, and the $f(\rho)_{-}$ branch asymptotic flat. In the cosmological case, the bounce allow to transit from $f(\rho)_{+}$ branch to $f(\rho)_{-}$ branch, whereas in the collapse scenario,it transitions from $f(\rho)_{-}$ branch to $f(\rho)_{+}$ branch. In both cases, the post-bounce branch describes the universe in which we currently live. Consequently, our investigation focuses on the phenomenologically relevant post-bounce sector.
In the cosmological case, the vacuum exterior associated with the post-bounce $f(\rho)_{-}$ branch contains a black hole, whereas in the collapse scenario, the vacuum exterior corresponding to the post-bounce $f(\rho)_{+}$ branch features only a cosmological horizon. Therefore, in what follows, we focus our quasinormal-mode (QNM) analysis on the $f(\rho)_{-}$ branch. We systematically investigate how angular momentum $l$ and the mass $M$ influence on effective potential, QNMs, and the late-time tail under scalar perturbations. We find that when angular momentum $l$ increases, the peak of the effective potential increases, and the quasinormal frequencies become longer-lived and dominate the time-domain signal. When the mass of black hole $M$ increases, the peak of the effective potential decreases. The late-time tails of black holes with different masses are similar and follow an inverse power-law behavior. To gain a better understanding of the differences in the structure of spacetime for the Schwarzschild black hole, the black hole in the qOS model, and the black hole in the qOS model with Lorentz term, we compared their effective potential and found a small deformation near the even horizon. Furthermore, we compared their time-domain profiles. The profile of the black hole in the qOS model closely matches that of the Schwarzschild black hole, while the black hole in the qOS model with Lorentz term exhibits a slower exponential decay rate. All three configurations display universal inverse power-law tails behavior across all mass configurations.
@@
The cosmological constant is generally believed to be the simplest explanation for the accelerated expansion of our Universe, while a negative cosmological constant introduces rich and distinctive physical signatures. Therefore, we consider the vacuum exterior metrics with a negative cosmological constant for both the $f(\rho)_{-}$ and $f(\rho)_{+}$ branches, and investigate their thermodynamic properties.
For the vacuum exterior metric corresponding to the $f(\rho)_{-}$ branch with a negative cosmological constant, we find that for small $r_{h}$, unlike the classical case where the Hawking temperature decreases monotonically from infinity, the temperature in the LQG-inspired black hole initially increases from zero, reaches a maximum, and then decreases. As $r_{h}$ becomes large, the behavior of the LQG-inspired black hole becomes similar to the classical scenario. For the heat capacity, we find that the LQG-inspired black hole exhibits an additional phase transition at a small $r_{h}$ compared to the classical scenario. When $r_{h}$ becomes large, the behavior of the LQG-inspired black hole coincides with that of the classical case. We also investigate the entropy of the LQG-inspired black hole. Compared to the classical case, the entropy-to-area ratio $\frac{S}{A}$ of the LQG-inspired black hole initially increases with $r_{h}$, reaches a maximum, and then decreases, eventually converging to the classical case at large $r_{h}$. This indicates that besides the classical area law $S=\frac{A}{4}$ term, the LQG-inspired black hole also has an additional logarithmic correction term due to the quantum gravity effect.
For the vacuum exterior metric corresponding to the $f(\rho)_{+}$ branch, we first consider the spacetime without a cosmological constant and compare its thermodynamic behavior with that of the classical Schwarzschild case. Since the quantum geometric effects in this model are equivalent to a positive effective cosmological constant $\Lambda_{eff}$, this comparison allows us to examine how such effects influence the thermodynamic properties. We find that for the LQG-inspired black hole, the Hawking temperature first decreases with increasing horizon radius $r_{c}$, and then increases linearly, while the Hawking temperature of de Sitter Schwarzschild black hole monotonically increases with $r_{c}$. However, the rate of increase differs between the two cases. The heat capacity of the LQG-inspired black hole exhibits an additional phase transition at small $r_{h}$ compared to the de Sitter Schwarzschild black hole. And at large horizon radius $r_{c}$, the heat capacity of the LQG-inspired black hole becomes negative, indicating that both black holes are thermodynamically unstable. Furthermore, as discussed above, even when a negative cosmological constant is introduced into the metric, provided that its absolute value does not exceed the effective cosmological constant $\Lambda_{eff}$, our main conclusions remain unchanged.